What is kinetic energy of relativistic particle.
1
Expert's answer
2020-09-10T14:40:06-0400

Total energy of the particle consists of rest energy and kinetic energy. The total relativistic energy of the particle is

"E = \gamma mc^2"

where "\gamma" is Lorenz factor.

The rest energy is "E_0 = mc^2". So kinetic energy is

"E_k = E - E_0 = \gamma m c^2 - mc^2 = mc^2 (\gamma -1)"